OpenStudy (anonymous):

NEED HELP The equation of a circle whose center is at (4, 0) and radius is length 2√(3) is (x - 4)² + y² = 2√3 (x - 4)² + y² = 12 (x + 4)² + y² = 12

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] replace \(h\) by \(4\) and \(k\) by \(0\) and \(r\) by \(2\sqrt{3}\)

OpenStudy (anonymous):